OpenStudy (anonymous):

what is the critical number? f(x) = 19 xln x

OpenStudy (anonymous):

my friend, that will take you quite a while

OpenStudy (anonymous):

\[f'(x)=19\ln(x)+19\]

OpenStudy (anonymous):

set \[f'(x)=0\] get \[\ln(x)+1=0\] so \[\ln(x)=-1\implies x = e^{-1}=\frac{1}{e}\]

OpenStudy (anonymous):

f(x) = 19 xln x f'(x) = 19 ( lnx + x/x) = 19 ( lnx + 1) f'(x ) = 0 --> lnx = -1 => x = e^-1 = 1/e